OpenStudy (anonymous):

Multiply and simplify if possible. 4^√11 4^√10

OpenStudy (anonymous):

It's hard to know what you mean if you don't use the equation editor. Please restate the question using the editor.

OpenStudy (anonymous):

4^√11 x 4^√10 = 11^(1/4) * 10^(1/4) = (110)^(1/4) = ^4√110 Or = ⁴√110 or, because they are the same, namely root root 4, can be multiplied directly = ^4√(11*10) = ⁴√110 hope i help

OpenStudy (anonymous):

@blackprincesssen Thanks, that cleared it up c:

OpenStudy (anonymous):

lol your welcome glade to help :D